OpenStudy (anonymous):

how do you graph: 5x - y greater than or equal to 5 and y < 4

OpenStudy (gorv):

5x-y>=5 5x-5>=y y<=5x-5 for x=0 y<=-5 for y=0 0<=5x-5 5<=5x x=>1